A soccer ball at rest is kicked by a soccer player. The ball travels a distance of 0.4 m over a time of 0.16 s. What is the acceleration rate of the ball over the period?


1
Expert's answer
2021-08-27T14:39:20-0400

"s=at^2/2 \to a=2s/t^2=2\cdot0.4/0.16^2=31.25\ (m/s^2)"